OpenStudy (trojanpoem):

Integrate:

OpenStudy (trojanpoem):

\[\int\limits_{0}^{1} e^{-y^2} dx\]

OpenStudy (michele_laino):

how is defined the relation between \(y\) and \(x\) ?

OpenStudy (trojanpoem):

No relation, the problem presents in Integration Kokeboken PDF.

OpenStudy (*poker-face*):

noob

OpenStudy (anonymous):

no relation between x and y so treat y as constant \(\int\limits_{0}^{1} x^0 e^{-y^2} dx\) \(\left[ x e^{-y^2} \right]^{1}_0\) =\(e^{-y^2}\)

OpenStudy (trojanpoem):

@bombdiggity, convincing.

OpenStudy (*poker-face*):

@bombdiggity ily<3